Find an equation in standard form of the parabola passing through the following points (-1,1) (0,5) (2,7)

Thank You.

Answer by lwsshak3(11628)

Find an equation in standard form of the parabola passing through the following points (-1,1) (0,5) (2,7)
y=Ax^2+Bx+C
(1) 1=A-B+C (-1,1)
(2) 5=C (0,5)
(3) 7=4A+2B+C (2,7)
..
1=A-B+5
7=4A+2B+5
2=2A-2B+10
add
9=6A+15
6A=-6
A=-1
..
B=A+5-1=3
Equation:
y=Ax^2+Bx+C
y=-x^2+3x+5
complete the square
y=-(x^2-3x+9/4)+9/4+5
y=-(x-3/2)^2+29/4
This is an equation of a parabola that opens downwards with vertex at(3/2,29/4)